OpenStudy (anonymous):

12-3(2w+1)=11w-3(11+w) solve the equation

OpenStudy (anonymous):

distribute: 12-6w-3=11w-33-3w simplify 9-6w=8w-33 +33 +6w 42=14w w= 3

OpenStudy (anonymous):

i think that's right

OpenStudy (anonymous):

one sec and ill let you know

OpenStudy (anonymous):

it was correct